# Contrastive Learning from Synthetic Audio Doppelgängers

### v1

@@ 3.5 Evaluation Tasks
To obtain a broad picture of the quality of our learned representations, we conduct experiments on a range of audio classification tasks selected from the HEAR [70] and ARCH [71] benchmarks. We are primarily interested in tasks that focus on real-world sounds (for example, vs. speech or music audio). We use a modified version of the HEAR evaluation setup to conduct linear probing experiments (instead of using an MLP probe) for a more direct signal of representation quality. We use the Adam optimizer to train this linear probe for the benchmark-specified number of epochs, with the default learning rate of 0.001 and a batch size of 32.
We evaluate on tasks covering a wide range of capabilities including sound classification tasks like ESC-50 [72], FSD-50k [73], and UrbanSound8K [74], vocal affect tasks with and without speech like VIVAE [75] and CREMA-D [76], musical pitch recognition via NSynth Pitch (5h) [27], vocal sound imitation recognition using Vocal Imitations [77], and LibriCount [78] for a "cocktail party" style speaker count estimation task.
@@ 4.1 Benchmark Results
In Table 1, we show results across 8 tasks. The top section features external baselines from the HEAR [70] leaderboard and ARCH [71] benchmark results, first the strongest overall and then only self-supervised. It also includes results from MS-CLAP [79] linear probing experiments, GURA [80] (strongest overall model on HEAR), and finally the original ResNet18 trained on VGGSound (supervised) [66]. Note that HEAR leaderboard results may use MLP probes, whereas ours are linear. We add additional internal baselines, including random weights, synthetic data trained without $\delta$ but with augmentations, and variants of ResNet18 we trained on VGGSound (with augmentations, and alternately with temporal jitter). Finally, we include a selection of our results; the best overall score we achieve using our synthetic approach (first row), followed by the best-performing model trained on data from each of the synthesizer architectures (including Voice with augmentations). In Section A.3, we provide a full set of results from all model variants: all synthetic datasets for all values of $\delta$, and further baselines less performant than those we present here.
@@
MS-CLAP Linear [79] 89.95 82.29 23.15 50.24 54.51
GURA (HEAR) [80] 74.35 38.20 75.21 41.32 18.48 68.34
VGGSound Sup. [66] 87.45 77.57 39.38 43.80 54.36 43.76 14.06 56.10
@@
Table 1: Evaluation results on a suite of tasks including (from left to right) ESC-50 [72], UrbanSound8k [74], VIVAE [75], NSynth Pitch 5h [27], CREMA-D [76], FSD50k [73], Vocal Imitation [77], and LibriCount [78]. For internal baselines, we only bold tasks where the baseline beats the best synthetically trained result. Results for all synthetic variants are in Section A.3.
@@ 4.2 Characterizing the Data Distribution
Here, we focus on understanding the distribution of synthetic sounds and how they differ from natural sound properties. We primarily use our alternate training set, VGGSound [66], for these measures. Unless specified otherwise, we use a randomly sampled (for VGGSound) or generated (for synthetic) set of 1000 sounds for each given dataset used for these characterizations. Our goal is to help understand what properties of the synthetic data make it useful for representation learning, given its strong performance.
@@ 4.2.1 Embedding Similarity
First, we look at the distribution of synthetic sound pairs and establish that $\delta$ meaningfully controls (a proxy for) perceptual or semantic dissimilarity. We operationalize this using LAION-CLAP embeddings [81], since they are trained on a large variety of sounds with semantic descriptions associated. Figure 1A shows how the average cosine similarity decreases monotonically with increasing $\delta$ for all 3 synthesizers. Figure 1B provides an alternate view of the space of positive pairs, compared to negatives (other samples). Here, we plot the first two principal components of the CLAP embeddings along with the path length for each positive pair of synthesized samples from a Voice synthesizer. As $\delta$ increases, the path lengths increase and overlap more resulting in less clear separation of positive pairs from negatives. We view this as a signal that we can effectively control the hardness of the contrastive task using $\delta$, the perturbation factor.
Figure 1: (A: Top) Average CLAP [81] embedding cosine similarity between positive pairs for different architectures and different values of $\delta$. (B: Bottom) PCA of CLAP embeddings for sounds generated with the Voice architecture, with line segments showing distances between paired examples. Red and blue points are paired positive instances. Across both plots, as $\delta$ increases, the positive pairs systematically become more perceptually dissimilar (via the CLAP embedding proxy).
@@ 4.2.2 Similarities and Differences from Real Data
Next, we compare the synthetic data distribution to VGGSound [66] data. Figure 2A compares a selection of features’ distributions between several dataset variants. For synthetic datasets, we have Voice, VoiceFM, and ParametricSynth variants. For real datasets, we have VGGSound. We first compare to randomly sampled 1-second chunks. Here, the synthetic sounds match several feature distributions well, such as Inharmonicity [82], Odd-to-Even Harmonic Ratio [83], Pitch Salience [84], and, to a lesser extent, Spectral Flatness [82]. However, the synthetic sounds have higher Spectral Flux [85] and Complexity [86]. Note that ParametricSynth also has lower pitch salience. We believe this is due to its larger mixture of sound generators which reduce salience of particular pitches.
@@
Figure 2: Comparisons of synthetic and real sound data (VGGSound [66]) on (A: Top) spectral features and (B: Bottom) causal uncertainty. Spectral features of synthetic sounds partially replicate real sounds, but exhibit differences in complexity and flux. Synthetic sounds are also more causally ambiguous, indicating a distribution shift. Using dense mixtures of real sounds partially closes these gaps, suggesting the synthetic sounds are different in part due to their density of auditory information.
@@ 4.2.3 Causal Uncertainty
We also consider causal uncertainty [87, 88, 89], a factor that we intuitively expect to be different for the synthetic sounds. Helmholtz famously discussed perception in terms of unconscious causal inference from sensory input [90], but the synthetic sounds have no physical causes and do not come from well-understood categories. In Figure 2B, we plot 3 proxies for causal uncertainty derived from probabilities of an AST classifier [91] trained on AudioSet. We use the formulation from prior work of $H_{cu}$, the maximum predicted probability [88, 89]. We also propose two simple metrics to corroborate this: $H_{p}$ the (normalized) entropy of the output probability distribution, and a confidence score $S_{conf}$, the difference in probability between the most and second-most probable classes (log-scaled for the plot). Across all, the synthetic sounds are more causally uncertain than the real sounds. However, as with the spectral feature distributions, using mixtures of VGGSound [66] clips moves the real distribution slightly closer to the synthetic distribution per $H_{cu}$ and $H_{p}$. We speculate that exposure to more causally uncertain sounds might be subtly helpful for representation learning; for example, they may contain diverse features that aid generalization to more ambiguous sounds present in downstream tasks. We characterize this as another important distributional difference between the synthetic sounds and realistic sounds from datasets such as VGGSound.
@@ 4.2.4 Similarity to Target Distributions
Another lens we can use to understand the effectiveness of training on synthetic data is in terms of the distribution of sounds in the target downstream tasks. A common metric to compare sound distributions is the Fréchet Audio Distance (FAD) [92]. For simplicity, we use the canonical formulation based on VGGish embeddings, though there are some limitations of this [93, 94], and we use either the validation sets or first multi-fold splits of the target task audio. Table 2 shows that for ESC-50 [72], VGGSound is closer in distribution to the target sounds, likely due to ESC-50’s focus on environmental sounds. For all other tasks, however, the synthetic sounds achieve a lower FAD, suggesting they may better capture task-relevant features for these tasks’ sounds. This finding echoes a study of MMDs in torchsynth [57], where the Voice architecture shows higher-than-expected similarity to FSD50k sounds. We hypothesize that the synthetic training allows the model to see a wide variety of spectral behavior rapidly, in a way that supports an array of tasks.
@@
Table 2: FAD [92] scores between different synthetic/real datasets and target downstream task data distributions, computed using either validation sets or the first fold (for multi-fold datasets). For 5/6 tasks, Voice achieves the lowest FAD despite containing synthetic sounds. On ESC-50, however, the VGGSound distribution appears to be closest.
@@ 4.3 Ablations and Sensitivity Analysis
In Figure 3, we study the effect of the perturbation factor $\delta$ on downstream task performance across all tasks for the strongest (Voice) architecture with and without additional (FX chain based) augmentations. Overall, $\delta=0.25$ appears to give the best result across tasks, with a few notable exceptions like NSynth without augmentations, and LibriCount overall.
@@ 5 Limitations
We opted for small-scale, established architectures like ResNets due to computational constraints and the desire for generalizable results. Scaling up to larger architectures like AST [91] could further validate our approach in high-compute environments (our attempts to train such models were infeasible). Additionally, we focused on direct comparisons between synthetic and real data, excluding hybrid approaches that might offer valuable insights, since the design space of hybrid approaches (e.g. mixtures or fine-tuning) is large and warrants a deeper exploration.
Our simple isometric Gaussian noise perturbation proved effective, but it relies on a simplifying independence assumption (parameters are correlated within components and component chains) and doesn’t account for parameter semantics (e.g. octaves in pitch). Future work could explore fine-grained, semantically-aware perturbations operationalized via the covariance structure of the perturbations. Scheduling or adapting the perturbation factor during training could offer more challenging examples. It could also be learned, which would require stable end-to-end auto-differentiation (challenging depending on the synthesizer architecture). Our evaluation focused on standard classification benchmarks, but incorporating aspects like representation disentanglement could offer a richer view. Finally, while we explored various synthesis options, the possibilities are vast, and future work could investigate variable architectures for enhanced data augmentation.
On a broader note, we believe it’s important to examine synthetic data-generating procedures for possible biases in the way that we do with datasets. Though we think this procedure can mitigate some of the biases in collectible real datasets, different synthesizer architectures, values of $\delta$, and other decisions might inadvertently produce performance gaps for different tasks, applications, and downstream populations of use. We evaluated on a wide range of tasks in part to explore this possibility, but further evaluations would be helpful to assess these impacts.
@@ Acknowledgements
The authors acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center for providing resources that have contributed to the research results reported within this paper. Manuel was supported by a US-Spain Fulbright grant.
@@ A.1 Comparison of Different Architectures Across Tasks
Figure 4: Scores with a fixed $\delta=0.25$ and different synthesizer architectures for a suite of tasks including (from left to right) UrbanSound8k [74], ESC-50 [72], LibriCount [78], CREMA-D [76], VIVAE [75], NSynth Pitch 5h [27], FSD50k [73], and Vocal Imitation [77]
@@ A.2 Effects of Increasing Perturbation Factor $\delta$ on Training
We seek to understand how increasing $\delta$ impacts the training dynamics. Figure 5 shows the impact of different $\delta$ on the final validation value of the alignment and uniformity costs respectively. Alignment cost increases monotonically with $\delta$, which shows the increased difficulty of aligning increasingly distant pairs. Uniformity, interestingly, has an inverted-U-shaped relationship with $\delta$. This may be because as the model struggles to align positives with moderate noise driven variation, it incurs a cost in uniformity in order to do so (e.g. creating clusters). With large $\delta$, the amount of noise present is significant, alignment is difficult, and the representations may be able to be more spread out.
@@ A.3 Results for all Variants
MS-CLAP Linear [79, 95] 89.95 82.29 23.15 50.24 54.51
GURA (HEAR) [80] 74.35 38.20 75.21 41.32 18.48 68.34
VGGSound Sup. [66] 87.45 77.57 39.38 43.80 54.36 43.76 14.06 56.10

### v2

@@ 3.5 Evaluation Tasks
To obtain a broad picture of the quality of our learned representations, we conduct experiments on a range of audio classification tasks from the HEAR (Turian et al., 2022) and ARCH (La Quatra et al., 2024) benchmarks. We use evaluation tasks that focus on general audio understanding, rather than tasks that are highly specialized or domain-specific (e.g., tasks exclusively related to speech or music). Our method aims to learn general-purpose audio representations from synthetic data; therefore, we implement tasks which encompass a broad range of everyday sounds. This aligns with our goal of demonstrating the effectiveness of our representations in diverse real-world scenarios.
These tasks cover a wide range of capabilities including sound classification tasks like ESC-50 (Piczak, 2015), FSD-50k (Fonseca et al., 2021a), and UrbanSound8K (Salamon et al., 2014), vocal affect tasks with and without speech like VIVAE (Holz et al., 2022) and CREMA-D (Cao et al., 2014), musical pitch recognition via NSynth Pitch (5h) (Engel et al., 2017), vocal sound imitation recognition using Vocal Imitations (Kim et al., 2018), and LibriCount (Stöter et al., 2018) for a “cocktail party” style speaker count estimation task. We conduct linear probing experiments using the Adam optimizer for the benchmark-specified epochs with the default learning rate of 0.001 and a batch size of 32.
@@ 4.1 Benchmark Results
In Table 1, we show results across 8 tasks. The top section features external baselines from the HEAR (Turian et al., 2022) leaderboard and ARCH (La Quatra et al., 2024) benchmark results, first the strongest overall and then only self-supervised. It also includes results from MS-CLAP (Elizalde et al., 2023) linear probing experiments, GURA (Wu et al., 2022) (strongest overall model on HEAR), and finally the original ResNet18 trained on VGGSound (supervised) (Chen et al., 2020a). Note that HEAR leaderboard results may use MLP probes, whereas ours are linear. We add additional internal baselines, including random weights, synthetic data trained without $\delta$ but with augmentations, and variants of ResNet18 we trained on VGGSound (with augmentations, and alternately with temporal jitter). Finally, we include a selection of our results; the best overall score we achieve using our synthetic approach (first row), followed by the best-performing model trained on data from each of the synthesizer architectures (including Voice with augmentations). In Section A.3, we provide a full set of results from all model variants: all synthetic datasets for all values of $\delta$, and further baselines less performant than those we present here.
@@
MS-CLAP Linear 89.95 82.29 23.15 50.24 54.51
GURA (HEAR) 74.35 38.20 75.21 41.32 18.48 68.34
VGGSound Sup. 87.45 77.57 39.38 43.80 54.36 43.76 14.06 56.10
@@
Table 1: Evaluation results on a suite of tasks including (from left to right) ESC-50 (Piczak, 2015), UrbanSound8k (Salamon et al., 2014), VIVAE (Holz et al., 2022), NSynth Pitch 5h (Engel et al., 2017), CREMA-D (Cao et al., 2014), FSD50k (Fonseca et al., 2021a), Vocal Imitation (Kim et al., 2018), and LibriCount (Stöter et al., 2018). For internal baselines, we only bold tasks where the baseline beats the best synthetically trained result. Results for all synthetic variants are in Section A.3.
@@ 4.2 Characterizing the Data Distribution
Here, we focus on understanding the distribution of synthetic sounds and how they differ from natural sound properties. We primarily use our alternate training set, VGGSound (Chen et al., 2020a), for these measures. Unless specified otherwise, we use a randomly sampled (for VGGSound) or generated (for synthetic) set of 1000 sounds for each given dataset used for these characterizations. Our goal is to help understand what properties of the synthetic data make it useful for representation learning, given its strong performance.
@@ 4.2.1 Embedding Similarity
First, we look at the distribution of synthetic sound pairs and establish that $\delta$ meaningfully controls (a proxy for) perceptual or semantic dissimilarity. We operationalize this using LAION-CLAP embeddings (Wu et al., 2023), since they are trained on a large variety of sounds with semantic descriptions associated. Figure 1A shows how the average cosine similarity decreases monotonically with increasing $\delta$ for all 3 synthesizers. Figure 1B provides a view of how $\delta$ affects the geometry of the embedding space. Here, we plot the first two principal components of the CLAP embeddings along with the path length for each positive pair of synthesized samples from a Voice synthesizer. As $\delta$ increases, the path lengths increase and overlap more resulting in less clear separation of positive pairs from negatives. We view this as a signal that we can effectively control the hardness of the contrastive task using $\delta$, the perturbation factor.
Figure 1: (A: Top) Average CLAP (Wu et al., 2023) embedding cosine similarity between positive pairs for different architectures and different values of $\delta$. (B: Bottom) PCA of CLAP embeddings for sounds generated with the Voice architecture, with line segments showing distances between paired examples. Red and blue points are paired positive instances. Across both plots, as $\delta$ increases, the positive pairs systematically become more perceptually dissimilar (via the CLAP embedding proxy).
@@ 4.2.2 Similarities and Differences from Real Data
Next, we compare the synthetic data distribution to VGGSound (Chen et al., 2020a) data. Figure 2A compares a selection of features’ distributions between several dataset variants. For synthetic datasets, we have Voice, VoiceFM, and ParametricSynth variants. For real datasets, we have VGGSound. We first compare to randomly sampled 1-second chunks. Here, the synthetic sounds match several feature distributions well, such as Inharmonicity (Peeters, 2004), Odd-to-Even Harmonic Ratio (Martin & Kim, 1998), Pitch Salience (Ricard, 2004), and, to a lesser extent, Spectral Flatness (Peeters, 2004). However, the synthetic sounds have higher Spectral Flux (Tzanetakis & Cook, 1999) and Complexity (Laurier et al., 2010). Note that ParametricSynth also has lower pitch salience. We believe this is due to its larger mixture of sound generators which reduce salience of particular pitches.
@@
Figure 2: Comparisons of synthetic and real sound data (VGGSound (Chen et al., 2020a)) on (A: Top) spectral features and (B: Bottom) causal uncertainty. Spectral features of synthetic sounds partially replicate real sounds, but exhibit differences in complexity and flux. Synthetic sounds are also more causally ambiguous, indicating a distribution shift. Using dense mixtures of real sounds partially closes these gaps, suggesting the synthetic sounds are different in part due to their density of auditory information.
@@ 4.2.3 Causal Uncertainty
We also consider causal uncertainty (Ballas & Sliwinski, 1986; Ananthabhotla et al., 2019; Boger et al., 2021), a factor that we intuitively expect to be different for the synthetic sounds. Helmholtz famously discussed perception in terms of unconscious causal inference from sensory input (Helmholtz, 1867), but the synthetic sounds have no physical causes and do not come from well-understood categories. In Figure 2B, we plot 3 proxies for causal uncertainty derived from probabilities of an AST classifier (Gong et al., 2021) trained on AudioSet. We use the formulation from prior work of $H_{cu}$, the maximum predicted probability (Ananthabhotla et al., 2019; Boger et al., 2021). We also propose two simple metrics to corroborate this: $H_{p}$ the (normalized) entropy of the output probability distribution, and a confidence score $S_{conf}$, the difference in probability between the most and second-most probable classes (log-scaled for the plot). Across all, the synthetic sounds are more causally uncertain than the real sounds. However, as with the spectral feature distributions, using mixtures of VGGSound (Chen et al., 2020a) clips moves the real distribution slightly closer to the synthetic distribution per $H_{cu}$ and $H_{p}$. We speculate that exposure to more causally uncertain sounds might be subtly helpful for representation learning; for example, they may contain diverse features that aid generalization to more ambiguous sounds present in downstream tasks. We characterize this as another important distributional difference between the synthetic sounds and realistic sounds from datasets such as VGGSound.
@@ 4.2.4 Similarity to Target Distributions
Another lens we can use to understand the effectiveness of training on synthetic data is in terms of the distribution of sounds in the target downstream tasks. A common metric to compare sound distributions is the Fréchet Audio Distance (FAD) (Kilgour et al., 2018). For simplicity, we use the canonical formulation based on VGGish embeddings, though there are some limitations of this (Gui et al., 2024; Tailleur et al., 2024), and we use either the validation sets or first multi-fold splits of the target task audio. Table 2 shows that for ESC-50 (Piczak, 2015), VGGSound is closer in distribution to the target sounds, likely due to ESC-50’s focus on environmental sounds. For all other tasks, however, the synthetic sounds achieve a lower FAD, suggesting they may better capture task-relevant features for these tasks’ sounds. This finding echoes a study of MMDs in torchsynth (Turian et al., 2021), where the Voice architecture shows higher-than-expected similarity to FSD50k sounds. We hypothesize that the synthetic training allows the model to see a wide variety of spectral behavior rapidly, in a way that supports an array of tasks.
@@
Table 2: FAD (Kilgour et al., 2018) scores between different synthetic/real datasets and target downstream task data distributions, computed using either validation sets or the first fold (for multi-fold datasets). For 5/6 tasks, Voice achieves the lowest FAD despite containing synthetic sounds. On ESC-50, however, the VGGSound distribution appears to be closest.
@@ 4.3 Ablations and Sensitivity Analysis
In Figure 3, we study the effect of the perturbation factor $\delta$ on downstream task performance across all tasks for the strongest (Voice) architecture with and without additional (FX chain based) augmentations. Overall, we found in early experiments that $\delta=0.25$ gives the best results. This is observed on the full suite of tasks (with notable exceptions like NSynth without augmentations, and LibriCount overall). In Section A.2, we discuss why this value might be strongest from the perspective of alignment and uniformity results.
@@ 5 Limitations
Our study demonstrates the efficacy of our approach using established architectures like ResNets, balancing computational efficiency with the goal of producing generalizable results. While we focused on these architectures, our findings lay the groundwork for future investigations with larger encoders such as AST (Gong et al., 2021); this is a straightforward extension. Our research also focused on a clear comparison between synthetic and real data, which allowed us to rigorously evaluate our method’s effectiveness. The potential for hybrid approaches, combining synthetic and real data, has a wide range of possibilities in mixing strategies and fine-tuning techniques. These can all be explored without changing our method itself.
The isometric Gaussian noise perturbation proved highly effective, despite its simplicity, since it changes identifiable attributes (e.g. pitch) subtly. This success points to the robustness of our method, while also highlighting opportunities for more sophisticated perturbation strategies to further enhance it. Future work could explore anisotropic perturbations that account for parameter relationships. Adaptive or learned perturbation strategies could also offer significant advancements. Additionally, our evaluation centered on widely used classification benchmarks to create a foundational assessment of the method’s performance. Expanding this evaluation to include metrics such as representation disentanglement could offer additional insights into the quality and utility of the synthetic data.
On a broader note, we believe it’s important to examine synthetic data-generating procedures for possible biases, similar to the scrutiny applied to real datasets. Though we think this procedure can mitigate some of the biases in real datasets, different synthesizer architectures, values of $\delta$, and other decisions might inadvertently produce performance gaps for different tasks, applications, and downstream populations of use. We evaluated on a wide range of tasks in part to explore this possibility, but further evaluations would be helpful to assess these impacts.
@@ Acknowledgements
Manuel received the support of a fellowship from “la Caixa” Foundation (ID 100010434). The fellowship code is LCF/BQ/EU23/12010079. The authors acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center for providing resources that have contributed to the research results reported within this paper. We extend our heartfelt thanks to all participants in the listening study.
@@ A.1 Comparison of Different Architectures Across Tasks
Figure 4: Scores with a fixed $\delta=0.25$ and different synthesizer architectures for a suite of tasks including (from left to right) UrbanSound8k (Salamon et al., 2014), ESC-50 (Piczak, 2015), LibriCount (Stöter et al., 2018), CREMA-D (Cao et al., 2014), VIVAE (Holz et al., 2022), NSynth Pitch 5h (Engel et al., 2017), FSD50k (Fonseca et al., 2021a), and Vocal Imitation (Kim et al., 2018)
@@ A.2 Effects of Increasing Perturbation Factor $\delta$ on Training
We seek to understand how increasing $\delta$ impacts the training dynamics. In particular, the alignment and uniformity objectives are in tension (Wang & Isola, 2020). A small $\delta$ leads to easy positive pairs (high similarity), resulting in low alignment cost but potentially poor generalization. Conversely, too large $\delta$ produces hard positive pairs (low similarity), increasing the alignment cost but potentially hindering optimization. The optimal $\delta$ should balance this trade-off, however the complexity of the synthesizer function and the embedding function make deriving a closed-form solution for this infeasible. As such, we must explore the effect empirically.
@@
Figure 5 shows the impact of different $\delta$ on the final validation value of the alignment and uniformity costs respectively. Alignment cost increases monotonically with $\delta$, which shows the increased difficulty of aligning increasingly distant pairs. Uniformity has an inverted-U-shaped relationship with $\delta$, suggesting that as the model struggles to align positives with moderate noise driven variation, it incurs a cost in uniformity in order to do so (e.g. creating clusters). With large $\delta$, the amount of noise present is significant, alignment is difficult, and the representations can be more spread out. The theoretically optimal value of $L_{unif}$ is $-2t=4$, which all values of $\delta$ remain close to. In Figure 8 of Wang & Isola (2020), the best performance with a more complex task and encoder (review classification) is observed when alignment is on the higher side (but not the maximum), and uniform is low (close to the optimal value). In our experiments, $\delta=0.25$ gets closest to this, and we observe it to be the strongest as well.
@@ A.3 Results for all Variants
MS-CLAP Linear 89.95 82.29 23.15 50.24 54.51
GURA (HEAR) 74.35 38.20 75.21 41.32 18.48 68.34
VGGSound Sup. 87.45 77.57 39.38 43.80 54.36 43.76 14.06 56.10